# Realistic Extreme Image Rescaling via Generative Latent Space Learning

### v1

@@ Abstract
Image rescaling aims to learn the optimal downscaled low-resolution (LR) image that can be accurately reconstructed to its original high-resolution (HR) counterpart. This process is crucial for efficient image processing and storage, especially in the era of ultra-high definition media. However, extreme downscaling factors pose significant challenges due to the highly ill-posed nature of the inverse upscaling process, causing existing methods to struggle in generating semantically plausible structures and perceptually rich textures. In this work, we propose a novel framework called Latent Space Based Image Rescaling (LSBIR) for extreme image rescaling tasks. LSBIR effectively leverages powerful natural image priors learned by a pre-trained text-to-image diffusion model to generate realistic HR images. The rescaling is performed in the latent space of a pre-trained image encoder and decoder, which offers better perceptual reconstruction quality due to its stronger sparsity and richer semantics. LSBIR adopts a two-stage training strategy. In the first stage, a pseudo-invertible encoder-decoder models the bidirectional mapping between the latent features of the HR image and the target-sized LR image. In the second stage, the reconstructed features from the first stage are refined by a pre-trained diffusion model to generate more faithful and visually pleasing details. Extensive experiments demonstrate the superiority of LSBIR over previous methods in both quantitative and qualitative evaluations. The code will be available at: https://github.com/wwangcece/LSBIR.
@@ 1 Introduction
With the explosive growth of ultra high-resolution digital visual data, image rescaling has become essential for enhancing data storage and transmission efficiency, optimizing resource usage in real-time applications, and ensuring proper display on devices with varying resolutions. During image rescaling, high-resolution (HR) images are often downscaled to visually pleasing low-resolution (LR) images for transmission or display. Subsequently, an upscaling process is employed to restore the downscaled images to their original resolution. Unlike the super-resolution (SR) task, image rescaling allows for joint optimization of both downscaling and upscaling processes, preserving detailed information crucial for effective upscaling.
Figure 1: Visual comparison of 16$\times$ image rescaling results. HCFlow struggles to obtain sufficient texture, while the reconstruction results of VQIR lack accurate semantic and structural information.
@@
In recent years, various image rescaling methods have been proposed, which can be classified into two main architectures: encoder-decoder [11, 22] and invertible neural networks [30, 16]. The encoder-decoder architecture employs two sets of neural networks to perform downscaling and upscaling operations. In contrast, invertible neural networks view image rescaling as a reversible process, using the forward and backward passes of an invertible neural network to simulate the downscaling and upscaling operations. Despite the significant improvements in reconstruction quality achieved by these methods, they are limited to small rescaling factors, such as 2$\times$ or 4$\times$.
@@
However, with the advancement of HR imaging devices, there is an increasing need for extreme rescaling factors (e.g., 16$\times$ or 32$\times$) to achieve more efficient storage and transmission of ultra HR visual data. When dealing with extreme rescaling factors, existing methods often struggle to generate sufficient details. Therefore, some approaches propose using pre-trained GANs as priors to constrain the solution space of the reconstructed results. GRAIN [36] and BDFlow [13] utilize pre-trained StyleGAN [8, 9] to generate high-quality upscaling face images. VQIR [29], on the other hand, leverages the high-quality visual embeddings encapsulated in pre-trained VQGAN [4] to achieve extreme rescaling on natural images. Although these methods have made significant progress in perceptual metrics, we find that their reconstruction results suffer from shortcomings in overall semantics and structure. As shown in Fig. 1, while VQIR generates richer details compared to HCFlow, its reconstruction results lack precise semantic and structural information. This is due to the use of both latent space and pixel space losses. The introduction of pixel-space loss can force the model to compromise on both optimization paths, leading to the generation of significant artifacts.
Figure 2: Image rescaling within pixel-space and latent-space. Different from methods that impose pixel loss, our approach applies loss only in the latent space, thereby ensuring the semantic integrity of the reconstruction results.
Therefore, we propose a novel framework named Latent Space Based Image Rescaling (LSBIR) for extreme rescaling, which does not impose any pixel loss on the reconstruction results, as illustrated in Fig. 2. Rescaling operations in the latent space have a broader receptive field, allowing richer information to be embedded into the LR images. Additionally, training in the latent space preserves the semantic integrity of the resampled features [7], better aligning with the intrinsic priors of pre-trained generative models. Specifically, LSBIR adopts a two-stage training strategy. In the first stage, a simple convolutional neural network (CNN) is used to rescale the latent code of the HR images. Concurrently, an invertible neural network (INN) completes the bidirectional mapping between the feature domain and the pixel domain. In the second stage, the rescaled features from the first stage guide the generation process of the pre-trained Stable Diffusion (SD) model, enhancing the details of the rescaled features. The first stage ensures that the rescaled features retain complete semantic information of the HR images, while the second stage generates richer textures and improves perceptual quality, balancing fidelity and realism in the final results.
@@
  • We propose a Latent Space Based Image Rescaling framework for extreme rescaling. Without imposing any pixel loss on the reconstruction results, our method exhibits excellent performance in global semantic and structural reconstruction.
  • We propose a novel decoupled strategy for feature rescaling. By decoupling the feature-pixel transform operation from the rescaling operation, this module can enhance the fidelity of feature reconstruction.
  • To the best of our knowledge, we are the first to apply diffusion models to the field of extreme image rescaling. We demonstrate that the rich knowledge contained in the pre-trained generative diffusion model can significantly enhance the perceptual quality of rescaling results.
Figure 3: Overview of proposed Latent-Space Based Image Rescaling (LSBIR). In the first training stage, we train a feature rescaling network in the latent space of a pre-trained VAE, which outputs LR images and rescaled features. In the second training stage, we use the features from the first stage as conditions and train an adapter to guide the generation process of the pre-trained SD model.
@@ 2.1 Image Rescaling
Image rescaling aims to downscale HR images to visually coherent LR images and then plausibly reconstruct the original HR images. Unlike image SR, image rescaling involves joint modeling of both downscaling and upscaling processes, leading to higher reconstruction accuracy. Earlier works, such as those by [11, 15], utilize CNNs to build an encoder-decoder framework, treating image rescaling as a unified task by jointly training the downscaling and upscaling processes. Instead of directly generating LR images, CAR [22] proposes generating content-adaptive resampling kernels based on HR images to perform the downsampling operation, thus avoiding pixel-level constraints on the downscaled images. To fully leverage the inherent reversibility of image rescaling, there has been a surge in the use of INNs for this task in recent years [14, 37, 33, 31]. IRN [30] is the first invertible framework that models image rescaling as a bijective transformation, embedding residual high-frequency components into a case-agnostic latent distribution for efficient reconstruction. HCFlow [16] assumes that high-frequency (HF) information depends on the LR image, thereby achieving better performance by incorporating LR conditions in HF information estimation. In the realm of extreme image rescaling, generative priors are usually employed to constrain the outcomes. Notably, GRAIN [36] and VQIR [29] leverage pre-trained StyleGAN [8, 9] and VQGAN [4], respectively, to achieve notable enhancements in perceptual quality. Nonetheless, these approaches are often limited in scope, either confined to handling images from specific domains or encountering difficulties when reconstructing faces and textual elements.
Recently, the emergence of the pre-trained SD model [20] has showcased its efficacy in both content generation [32] and image restoration [25]. Despite this progress, to date, no effort has been made to integrate it within the framework of extreme image rescaling, presenting an unexplored avenue with potential for advancing the field.
@@ 2.2 Diffusion Model for Image SR
Here, we briefly introduce the application of diffusion models in a related task, i.e., image SR. Diffusion models promote the development of natural image SR in two main ways: training diffusion models from scratch and leveraging pre-trained SD model. The former is represented by SR3 [21], which achieves perceptual quality comparable to GANs. The latter is represented by StableSR [25], which aims to build a model capable of handling image restoration under any complex degradation in real-world scenarios.
However, due to the inherent difficulty of integrating diffusion models into existing encoder-decoder architectures, few work has yet attempted to combine them with extreme image rescaling. In this work, we decompose extremec image rescaling into two sub-tasks: feature rescaling in the latent space and perception-oriented optimization. By doing so, we could explore the potential application of diffusion models in this field.
@@ 3 Methodology
3.1 Overview of LSBIR
Our LSBIR architecture, shown in Fig. 3, involves two training stages: Feature Rescaling in the Latent Space and Perception-oriented Optimization. Initially, the input HR image is encoded into a latent feature $z$, which is then rescaled by the Decoupled Feature Rescaling Module (DFRM) to generate an LR image $y$ and a rescaled latent feature $\hat{z}$. The first stage aims to align $\hat{z}$ with $z$ while ensuring $y$ retains the content of the original HR image, preserving semantic information for the second stageś refinement. Although decoding $\hat{z}$ produces high-fidelity HR images, they lack texture detail. Thus, in the second stage, we use an adapter to guide the pre-trained SD model, generating more realistic textures. During inference, we could balance fidelity and realism by weighting $\hat{z}$ from both stages.
@@ 3.2 Feature Rescaling in the Latent Space
3.2.1 Overall Network Architecture
In the first stage, we perform rescaling of the HR features $z$ in the latent space, obtaining LR images $y$ and rescaled features $\hat{z}$. A straightforward idea is to construct a transformation chain as $y=G_{e}\left(z\right)$, $\hat{z}=G_{d}\left(y\right)$. Here, $G_{e}$ is an encoder used to downscale the input features to the target size, while $G_{d}$ represents the decoder used to reconstruct the original high-resolution features. Guidance loss $\mathcal{L}_{gui.}$ and reconstruction loss $\mathcal{L}_{rec.}$ are applied to $y$ and $\hat{z}$ respectively, to force $y$ to have the same content as $x$, and $\hat{z}$ to be as similar as possible to $z$:
$\mathcal{L}_{\mathrm{gui.}}=\parallel y-\mathrm{Bicubic}(x)\parallel_{p}$ (1)
$\mathcal{L}_{\mathrm{rec.}}=\parallel\hat{z}-z\parallel_{p}$ (2)
However, these two losses are contradictory to each other. When the guidance loss approaches zero, the rescaling problem degenerates into a classical SR [28] problem. Conversely, when the reconstruction loss reaches its global optimum, the rescaling problem degenerates into an image compression problem [18]. Therefore, directly applying the guidance loss $\mathcal{L}_{gui.}$ on $y$ might negatively impact subsequent upscaling.
Considering the above issues, we devise the DFRM as shown in Fig. 4. Inspired by FGRN [14], we use a set of CNNs to perform downscaling and upscaling, and a separate set of INNs to perform the bidirectional mapping between the feature domain and the pixel domain. To achieve balance between feature rescaling accuracy and visual quality of LR images, we design another pixel guidance module in the encoder to incorporate pixel-level information. Specifically, DFRM perform feature rescaling through two independent transformation chains $\left(\left(x,z\right)\rightarrow w\rightarrow\hat{z}\right)$; $\left(w\leftrightarrow y\right)$ as follows:
$w=G_{e}\left(x,z\right)$ (3)
$w\approx F^{-1}\left(F\left(w\right)\right)=F^{-1}\left(y\right)$ (4)
$\hat{z}=G_{d}\left(w\right)$ (5)
Here, $F$ and $F^{-1}$ represent the forward and reverse processes of the INN, respectively. Given the latent variable $z\in R^{H/8\times W/8\times 4}$ of the pre-trained SD VAE as input, the encoder $G_{e}$ first transforms it into a compact representation $w\in R^{H/N\times W/N\times 3}$, whose size is the same as that of the target LR image and $N$ represents the target downscaling factor. Subsequently, a set of INNs perform the forward process to transform $w$ from the feature domain to the pixel domain and obtain the LR image $y\in R^{H/N\times W/N\times 3}$ through quantization. Symmetrically, in the upscaling process, the inverse of the INN is used to process the LR image $y$ and reconstruct $w$. Finally, the decoder $G_{d}$ is utilized to obtain the rescaled feature $\hat{z}$.
@@ 3.2.1 Overall Network Architecture
3.2.2 Domain Converter
The domain converter module consists of a series of stacked invertible blocks [3], which can be defined as $F=f_{1}\circ f_{2}\circ...f_{K}$. For the $i$-th invertible block $f_{i}$ in the forward process, the input $w_{k}$ is split into $w_{k}^{1}$ and $w_{k}^{2}$ along the channel axis, and they undergo affine coupling layers for information fusion as follows:
$w_{k+1}^{1}=w_{k}^{1}+\phi\left(w_{k}^{2}\right)$ (6)
$w_{k+1}^{2}=w_{k}^{2}\otimes\exp\left(\rho\left(w_{k+1}^{1}\right)\right)+\eta\left(w_{k+1}^{1}\right)$ (7)
Here, $\phi\left(\cdot\right)$, $\rho\left(\cdot\right)$ and $\eta\left(\cdot\right)$ could be arbitrary functions and we adopt the widely used RRDB [26] for its good representation ability. Finally, the output of the last invertible module, i.e., $f_{K}$, is quantized to obtain the LR image $y$. Similarly, in the reverse process, the LR image $y$ will be used as an input to reconstruct $w$. For the $i$-th invertible block $f_{i}$, the input $w_{k+1}$ is split into $w_{k+1}^{1}$ and $w_{k+1}^{2}$ and undergo affine coupling layers as follows, thus obtaining the low-resolution representation $w$:
$w_{k}^{2}=\left(w_{k+1}^{2}-\eta\left(w_{k+1}^{1}\right)\right)\otimes\exp\left(-\rho\left(w_{k+1}^{1}\right)\right)$ (8)
$w_{k}^{1}=w_{k+1}^{1}-\phi\left(w_{k}^{2}\right)$ (9)
3.2.3 Loss Function
@@
First stage training involves two losses: reconstruction loss and guidance loss. Considering the information lost in quantization operations, the bidirectional mapping between $w$ and $y$ is not exactly reversible. To ensure the robustness of the network to quantization operations, we consider two types of transformation chains in designing the reconstruction loss: $\left(x,z\right)\rightarrow w\rightarrow\hat{z}$ and $\left(x,z\right)\rightarrow w\rightarrow y\rightarrow\hat{w}\rightarrow\hat{z}$, which can be expressed as the following formula:
$\begin{split}\mathcal{L}_{\mathrm{rec.}}=&\parallel G_{d}\left(G_{e}\left(x,z\right)\right)-z\parallel_{p}+\\ &\parallel G_{d}\left(F^{-1}\left(F\left(G_{e}\left(x,z\right)\right)\right)\right)-z\parallel_{p}\end{split}$ (10)
@@
This combination can facilitate the network to reconstruct $\hat{z}$ with high-fidelity under both the existence and absence of quantization operations. Additionally, considering that the guidance loss should be decoupled from the reconstruction loss as much as possible, we design it as follows:
$\mathcal{L}_{\mathrm{gui.}}=\parallel\mathrm{sg}\left(F\left(G_{e}\left(x,z\right)\right)\right)-\mathrm{Bicubic}(x)\parallel_{p}\$ (11)
Here, $\mathrm{sg}\left(\cdot\right)$ is the stop-gradient operation preventing the gradient flowing into $G_{e}$ or $G_{d}$, which decouples the feature-pixel domain bidirectional mapping and feature rescaling operations. Finally, the loss in the first stage is a combination of these two losses:
$\mathcal{L}_{\mathrm{stage1}}=\lambda_{1}\mathcal{L}_{\mathrm{rec.}}+\lambda_{2}\mathcal{L}_{\mathrm{gui.}}\$ (12)

### v2

# One-step Generative Diffusion for Realistic Extreme Image Rescaling

@@ Abstract
Image rescaling aims to learn the optimal low-resolution (LR) image that can be accurately reconstructed to its original high-resolution (HR) counterpart, providing an efficient image processing and storage method for ultra-high definition media. However, extreme downscaling factors pose significant challenges to the upscaling process due to its highly ill-posed nature, causing existing image rescaling methods to struggle in generating semantically correct structures and perceptual friendly textures. In this work, we propose a novel framework called One-Step Image Rescaling Diffusion (OSIRDiff) for extreme image rescaling, which performs rescaling operations in the latent space of a pre-trained autoencoder and effectively leverages powerful natural image priors learned by a pre-trained text-to-image diffusion model. Specifically, OSIRDiff adopts a pseudo-invertible module to establish the bidirectional mapping between the latent features of the HR image and the target-sized LR image. Then, the rescaled features are refined by a pre-trained diffusion model to generate more faithful and visually pleasing details. The entire model is end-to-end trained to enable the diffusion priors to guide the rescaling process. Considering the spatially non-uniform reconstruction quality of the rescaled latent features, we propose a novel time-step alignment strategy, which can adaptively determine the generative strength of the diffusion model based on the degree of latent reconstruction errors. Extensive experiments demonstrate the superiority of OSIRDiff over previous methods in both quantitative and qualitative evaluations. The code will be available at: https://github.com/xxx/xxx.
@@ 1 Introduction
With the explosive growth of ultra high-resolution (HR) images, image rescaling has become essential for enhancing data storage and transmission efficiency [44], optimizing resource usage in real-time applications [26], and ensuring proper display on devices with varying resolutions [25]. During image rescaling, HR images are downscaled to low-resolution (LR) images for storage, transmission, or display. When HR images are needed, an upscaling process is employed to restore the downscaled images to their original resolution. Unlike image super-resolution (SR) task, image rescaling allows for joint optimization of both downscaling and upscaling processes, preserving detailed information crucial for effective upscaling.
Figure 1: Visual comparison of 16$\times$ image rescaling results.
@@
Recently published image rescaling methods can be classified into two main categories: encoder-decoder architecture [14, 31] and invertible neural networks [41, 18]. The encoder-decoder architecture employs two neural networks to perform downscaling and upscaling. While, invertible neural networks view image rescaling as a reversible process, using the forward and backward passes of an invertible neural network to simulate the downscaling and upscaling. Despite the significant improvements in reconstruction quality achieved by these methods, they are limited to small rescaling factors, such as 2$\times$ or 4$\times$.
@@
However, with the advancement of ultra HR imaging devices, there is an increasing need for extreme rescaling factors (e.g., 16$\times$ or 32$\times$) to achieve more efficient storage and transmission of ultra HR data [53, 8]. Existing image rescaling methods often struggle to generate sufficient detail under extreme rescaling factors. Therefore, some approaches propose using pre-trained GANs as priors to constrain the solution space of the reconstructed results. GRAIN [53] and BDFlow [15] use pre-trained StyleGAN [10, 11] to generate high-quality upscaling face images. VQIR [39] leverages the high-quality visual embeddings encapsulated in pre-trained VQGAN [4] to achieve extreme rescaling on natural images. Although these methods have made significant progress in perceptual metrics, we find that their reconstruction results suffer from incorrect semantics and structure. As shown in Fig. 1, although VQIR generates richer details compared to HCFlow, it still fails to reconstruct accurate text and image patterns.
To address the aforementioned issues, we propose a novel framework named One-Step Image Rescaling Diffusion (OSIRDiff) for extreme image rescaling, which performs image rescaling in the latent space of a pre-trained autoencoder and leverages the natural image priors stored in large text-to-image diffusion models to restore HR images. Specifically, first, we use an autoencoder to rescale the latent code of the HR image to the target size, and employ a set of invertible neural networks (INN) to perform the mapping from the feature space to the pixel space to obtain the LR image. Then, the rescaled features are fed into a pre-trained diffusion model to perform a single denoising step, thereby enhancing their perceptual quality. Finally, the entire model is trained end-to-end to allow the diffusion priors to guide the rescaling process. Due to the spatially non-uniform reconstruction quality of the rescaled latent features, different regions of the image require varying levels of generative capability for restoration. Previous studies have shown that the time-step in diffusion models directly controls the generative capacity of the model [35]. Therefore, we propose a time-step alignment mechanism that predicts the appropriate time-step based on the reconstruction quality of the rescaled latent features, thereby achieving a balance between fidelity and realism. As shown in Fig. 1, our model significantly enhances the overall quality of the rescaled images compared to other methods.
@@
  • We propose a One-Step Image Rescaling Diffusion framework that performs well in global semantic and structural reconstruction for the extreme image rescaling task.
  • We propose a novel decoupled strategy for feature rescaling. It decouples the generation of LR images from the rescaling operation, thereby improving the reconstruction accuracy of the rescaling process.
  • We design a time-step alignment strategy to address the spatially non-uniform reconstruction quality of the rescaled latent features, allowing for more effective utilization of the prior information from the text-to-image diffusion model.
Figure 2: Overview of the proposed One-Step Image Rescaling Diffusion (OSIRDiff). First, the HR images $x$ are encoded to obtain the latent features $z$. These latent features are then rescaled to the target size using the decoupled feature rescaling module (DFRM), producing the LR images $y$ and outputting the rescaled latent features $\hat{z}$. Next, the rescaled latent features $\hat{z}$ are passed into a pre-trained diffusion model to perform a single denoising step for perceptual enhancement to obtain $\hat{z}_{0}$. Finally, the perceptually enhanced latent features $\hat{z}_{0}$ are decoded to obtain the rescaled image $\hat{x}$.
@@ 2.1 Image Rescaling
Image rescaling aims to downscale HR images to visually coherent LR images and then plausibly reconstruct the original HR images. Unlike image SR, image rescaling involves joint modeling of both downscaling and upscaling processes, leading to higher reconstruction accuracy. Earlier researches [14, 17] utilize CNNs to build an encoder-decoder framework, treating image rescaling as a unified task by jointly training the downscaling and upscaling processes. Instead of directly generating LR images, CAR [31] proposes generating content-adaptive resampling kernels based on HR images to perform the downsampling operation, avoiding pixel-level constraints on the downscaled images. To fully leverage the inherent reversibility of image rescaling, there has been a surge in the use of INNs for this task [16, 54, 48, 43]. IRN [41] is the first invertible framework that models image rescaling as a bijective transformation, embedding residual high-frequency (HF) components into a case-agnostic latent distribution for efficient reconstruction. HCFlow [18] assumes that the HF information depends on the LR image, thereby achieving better performance by incorporating LR conditions in HF information estimation. In extreme image rescaling, generative priors are usually employed to constrain the outcomes. GRAIN [53] and VQIR [39] leverage pre-trained StyleGAN [10, 11] and VQGAN [4], respectively, to achieve notable enhancements in perceptual quality. Nonetheless, these approaches are often limited in image contents, either confined to handling images from specific domains or encountering difficulties when reconstructing faces and textual elements.
Recently, the emergence of the pre-trained Stable Diffusion (SD) model [27] has showcased its efficacy in both content generation [47] and image restoration [34]. Despite this progress, to date, no effort has been made to investigate it within the framework of extreme image rescaling, presenting an unexplored avenue with potential for advancing the field.
@@ 2.2 Diffusion Model for Image SR
Here, we briefly introduce the application of diffusion models in a related task, i.e., image SR. Diffusion models promote the development of natural image SR in two main ways: training diffusion models from scratch and leveraging pre-trained SD model. The former is represented by SR3 [28], which achieves perceptual quality comparable to GANs. The latter is represented by StableSR [34], which aims to build a model capable of handling image restoration under any complex degradation in real-world scenarios. However, due to the inherent difficulty of integrating diffusion models into existing encoder-decoder architectures, few work has yet attempted to combine them with extreme image rescaling. In this work, we decompose extreme image rescaling into two sub-tasks: feature rescaling in the latent space and one-step perception enhancement with the pre-trained SD U-Net.
2.3 Diffusion Acceleration
To accelerate the inference process of diffusion models, researchers have proposed various optimization strategies. Fast diffusion samplers [30, 12, 21, 51] effectively reduce the sampling steps required by pre-trained diffusion models, lowering them from the initial thousand steps to as few as 20–50 steps. However, further reductions often lead to substantial performance degradation. Distillation techniques [23, 29, 42, 52, 45, 24] have shown significant promise in accelerating inference. For example, Progressive Distillation [29] minimizes sampling steps to under 10 by training a sequence of student models with progressively increasing stride. Instaflow [20] advances this further, achieving one-step generation using reflow and distillation methods. Recently, score-based distillation approaches, including Distribution Matching Distillation [45] and Variational Score Distillation [24], have utilized pre-trained text-to-image diffusion models as distribution matching losses to enable one-step text-to-image generation. Although these methods have made significant strides, high training costs remain a challenge. Inspired by Diffusion2GAN [9], we propose a low-overhead training approach based on Low-Rank Adaptation (LoRA) [6], achieving one-step acceleration for diffusion models and adaptability for image rescaling tasks.
@@ 3 Methodology
3.1 Overview of OSIRDiff
As illustrated in Fig. 2, OSIRDiff consists of four parts: Latent Encoding, Feature Rescaling in the Latent Space, One-Step Perception Enhancement and Latent Decoding. First, the input HR images $x$ are mapped to the latent features $z$ via the pre-trained VAE encoder $\mathcal{E}$. Subsequently, the proposed Decoupled Feature Rescaling Module (DFRM) rescales $z$ to the target size, outputting LR images $y$ and the rescaled latent features $\hat{z}$. Our aim is to make $\hat{z}$ as close as possible to $z$ so that the rescaled results align with the latent space of the pre-trained VAE, while ensuring that $y$ and the original HR images $x$ share the same content. Then, the rescaled features $\hat{z}$ are fed into a denoising U-Net network to perform a single denoising step to get the perceptually enhanced latent features $\hat{z}_{0}$. Finally, perceptually enhanced latent features $\hat{z}_{0}$ are decoded via the pre-trained VAE decoder $\mathcal{D}$ to obtain the rescaled image $\hat{x}$. To dynamically adjust the generative capability based on the reconstruction quality of the rescaled latent features, we design a time-step alignment strategy, which consists of a time-step prediction module (TPM) and a refiner module.
@@ 3.2 Feature Rescaling in the Latent Space
3.2.1 Network Architecture.
In this part, we perform rescaling of the HR latent features $z$, obtaining LR images $y$ and rescaled features $\hat{z}$. A straightforward idea is to construct a transformation chain as $y=G_{e}\left(z\right)$, $\hat{z}=G_{d}\left(y\right)$. Here, $G_{e}$ is the encoder used to downscale the input features to the target sized LR images, while $G_{d}$ represents the decoder used to reconstruct the original HR features. Guidance loss $\mathcal{L}_{gui.}$ and reconstruction loss $\mathcal{L}_{rec.}$ are applied to $y$ and $\hat{z}$ respectively, to force $y$ to have the same content as $x$, and $\hat{z}$ to be as similar as possible to $z$:
$\mathcal{L}_{\mathrm{gui.}}=\parallel y-\mathrm{Bicubic}(x)\parallel_{1}$ (1)
$\mathcal{L}_{\mathrm{rec.}}=\parallel\hat{z}-z\parallel_{1}$ (2)
However, these two losses are contradictory to each other. When the guidance loss approaches zero, the rescaling problem degenerates into a classical SR problem [38], and little information regarding the downscaling process can be embedded into the LR image. In contrast, when the reconstruction loss reaches its global optimum, the rescaling problem degenerates into an image compression problem [22], which results in the LR image lacking any visual content. Therefore, directly applying the guidance loss $\mathcal{L}_{gui.}$ on $y$ could adversely impact subsequent upscaling.
@@
Considering the above issues, we devise the DFRM as shown in Fig. 3, which divides the rescaling operation into two sub-tasks: feature rescaling and LR image generation. Inspired by FGRN [16], we use a set of CNNs to perform downscaling and upscaling, and a separate INN to perform the bidirectional mapping between the feature domain and the pixel domain. To achieve a balance between feature rescaling accuracy and visual quality of LR images, we introduce another pixel guidance module in the encoder to incorporate pixel-level information, which can improve the quality of the LR image. Specifically, DFRM perform feature rescaling through two independent transformation chains $\left(\left(x,z\right)\rightarrow z_{\text{lr}}\rightarrow\hat{z}\right)$; $\left(z_{\text{lr}}\leftrightarrow y\right)$ as follows:
$z_{\text{lr}}=G_{e}\left(x,z\right)$ (3)
$z_{\text{lr}}\approx F^{-1}\left(F\left(z_{\text{lr}}\right)\right)=F^{-1}\left(y\right)$ (4)
$\hat{z}=G_{d}\left(z_{\text{lr}}\right)$ (5)
Here, $F$ and $F^{-1}$ represent the forward and backward processes of the INN, respectively. Given the latent variable $z\in R^{\frac{H}{8}\times\frac{W}{8}\times 4}$ of the pre-trained SD VAE as input, the encoder $G_{e}$ first transforms it into a compact representation $z_{\text{lr}}\in R^{\frac{H}{N}\times\frac{W}{N}\times 3}$, whose size is the same as that of the target LR image and $N$ represents the target downscaling factor. Subsequently, a set of INNs perform the forward process to transform $z_{\text{lr}}$ from the feature domain to the pixel domain and obtain the LR image $y\in R^{\frac{H}{N}\times\frac{W}{N}\times 3}$ by quantization. Symmetrically, in the upscaling process, the inverse of the INN is used to process the LR image $y$ and reconstruct $z_{\text{lr}}$. Finally, the decoder $G_{d}$ is utilized to obtain the rescaled feature $\hat{z}$.
3.2.2 Loss Function.
@@
Training of this part involves two losses: reconstruction loss and guidance loss. Considering the information lost in quantization operations, the bidirectional mapping between $z_{\text{lr}}$ and $y$ is not exactly reversible. To ensure the robustness of the network to quantization operations, we consider two types of transformation chains in designing the reconstruction loss: $\left(x,z\right)\rightarrow z_{\text{lr}}\rightarrow\hat{z}$ and $\left(x,z\right)\rightarrow z_{\text{lr}}\rightarrow y\rightarrow\hat{y}_{lr}\rightarrow\hat{z}$, which can be expressed as the following formula:
$\begin{split}\mathcal{L}_{\mathrm{rec.}}=&\parallel G_{d}\left(G_{e}\left(x,z\right)\right)-z\parallel_{1}+\\ &\parallel G_{d}\left(F^{-1}\left(F\left(G_{e}\left(x,z\right)\right)\right)\right)-z\parallel_{1}\end{split}$ (6)
@@
This combination can facilitate the network to reconstruct $\hat{z}$ with high fidelity in both the existence and absence of quantization operations. Additionally, the guidance loss is designed as follows:
$\mathcal{L}_{\mathrm{gui.}}=\parallel F\left(G_{e}\left(x,z\right)\right)-\mathrm{Bicubic}(x)\parallel_{1}\$ (7)
Finally, the loss function of this part is a combination of these two losses:
$\mathcal{L}_{\mathrm{res.}}=\lambda_{\mathrm{rec.}}\mathcal{L}_{\mathrm{rec.}}+\lambda_{\mathrm{gui.}}\mathcal{L}_{\mathrm{gui.}}\$ (8)
Figure 4: The rescaling operation leads to spatially non-uniform reconstruction quality of the rescaled latent features, which depends on the content of the image and the rescaling factor.
3.3 One-Step Perception Enhancement
The latent variable $\hat{z}$ obtained in Sec. 3.2 retains sufficient semantic information about the HR image, but lacks adequate texture. Therefore, we are proposed to leverage the strong priors stored in the pre-trained SD model to enhance the perceptual quality of the rescaling results in this part. Considering the original diffusion model requires a multi-step iterative sampling algorithm, this significantly increases inference time and makes it difficult to scale to practical applications. Inspired by recent work on accelerating diffusion models [9], we adopt an one-step inference strategy. Specifically, we can directly feed the rescaled latent variable $\hat{z}$ into the denoising U-Net $\epsilon=\epsilon_{\theta}\left(\hat{z},t\right)$ to perform a single denoising step as follows:
$\hat{z}_{0}=\frac{\hat{z}-\sqrt{1-\bar{\alpha}_{t}}\epsilon}{\sqrt{\bar{\alpha}_{t}}}$ (9)
where $\hat{z}_{0}$ represents the perceptually enhanced latent features, $\bar{\alpha}_{t}$ is the noise scheduler of the forward diffusion process of the SD model, and $t$ represents the time-step associated with the noise intensity. Finally, $\hat{z}_{0}$ can be transformed to the pixel domain through the pre-trained VAE decoder $\mathcal{D}$ to obtain the upscaled image $\hat{x}$:
$\hat{x}=\mathcal{D}\left(\hat{z}_{0}\right)$ (10)
Then, we can adopt the loss function between $\hat{x}$ and $x$ to optimize the U-Net using LoRA layers.
It should be noted that one of the key designs is the selection of the time-step $t$. Previous studies [35] have shown that the time-step in diffusion models could determine the trade-off between realism and fidelity. A larger $t$ results in stronger diversity and greater realism, but at the cost of fidelity. However, we have observed that the quality of the rescaled latent features exhibit non-uniform characteristics, as shown in Fig. 4. After rescaling, the low-frequency regions of an image typically experience less degradation, while the degradation in high-frequency areas is more conspicuous, necessitating a higher generative capability for restoration. Thus, we propose to dynamically predict the time-step $t$ based on the quality of the rescaled latent features. We propose to use a simple yet effective network named time-step prediction module (TPM) to regress $t$ based on the rescaled latent features:
$t=\mathrm{TPM}\left(\hat{z}\right)$ (11)
Subsequently, the predicted time-step $t$ is used to perform the single denoising step. During the training phase, we compute a single time-step for each input image patch. In the inference phase, we adopt a tiled inference approach, allowing for predicting different time-steps for various regions of a single image. However, the fixed scheduler used in equation (9) is not differentiable with respect to $t$, which prevents gradients from being back-propagated to the TPM. Considering that given $\hat{z}$, $\epsilon$ and $t$, equation (9) essentially represents a linear transformation, which can be learned by a neural network. However, if we simply use a randomly initialized neural network to simulate this process, it would cause the output to deviate significantly from the latent space of the pre-trained VAE, leading to training instability. Therefore, we propose a hybrid refiner module consisting of both a fixed scheduler and a learnable scheduler. The former uses equation (9) to perform the denoising step, while the latter employs a neural network to simulate the denoising scheduler. To ensure training stability, the final layer of the learnable scheduler is a zero-initialized convolution, allowing it to gradually correct the results of the fixed scheduler as training progresses. This process can be described as follows:
$\begin{split}\hat{z}_{0}=\mathcal{S}_{\mathrm{fixed}}\left(\hat{z},\epsilon,t_{0}\right)+\mathcal{S}_{\mathrm{learned}}\left(\hat{z},\epsilon,t\right)\end{split}$ (12)
Here $t_{0}$ represents the fixed pre-set time-step.
Figure 5: The structure of the refiner module. It combines both the fixed scheduler and learnable scheduler to perform the denoising process. The two schedulers are connected by a zero-convolution layer to stabilize the training.